OpenStudy (anonymous):

Formic acid (HCOOH) is secreted by ants. Calculate [H3O+] for a 1.18E-2 M aqueous solution of formic acid (Ka = 1.80E-4).

OpenStudy (aaronq):

write an equilibrium expression for the dissociation of formic acid. Plug in your values and solve for \([H_3O^+]\).

OpenStudy (anonymous):

Which one would I use as the products/reactants?

OpenStudy (anonymous):

Would it be HCOOH <==> H30+HCOO?

OpenStudy (anonymous):

Giving me, [H30][HCOO]/[HCOOH]?

OpenStudy (aaronq):

yes, thats right.

OpenStudy (anonymous):

so 1.8E-4=[H30][HCOO]/1.18E-2 1.8E-4*1.18E-2=[H3O][HCOO] 2.124E-6=[H3O][HCOO] then what?

OpenStudy (aaronq):

\([H_3O^+]=[HCOO^-]=x\); because each dissociation yields 1 molecule of each. so \(2.124*10^{-6}=x^2\) solve for x

OpenStudy (anonymous):

\[\sqrt{2.124e-6}=x\] x=.001457??

OpenStudy (aaronq):

yep, thats it.